# Beyond Memorization: Gradient Projection Enables Selective Learning in Diffusion Models

### v1

@@ 5 Conclusions, Limitations, and Future Work
Addressing the computational overhead is a primary direction for future optimization, potentially through low-rank or sparse gradient estimation techniques. Concurrently, future research should focus on optimizing $\lambda$ to maintain training stability while achieving the theoretical guarantee of $\lambda=1$ (perfect excising). Further work is also needed on disentangling complex copyrighted attributes that remain implicitly entangled with desirable text-conditioned aspects of the image. Additionally, integrating gradient projection into video diffusion pipelines can ensure spatiotemporal copyright compliance without bespoke architectural changes, facilitating cross-modal research and encouraging future advancements in industry best practices for IP protection across the foundation model ecosystem.
6 Conclusions, Limitations, and Future Work
We formalized copyright preservation as a stringent security requirement, demanding the prevention of any identifiable proprietary attributes. Our gradient projection framework systematically removes training signals aligned with protected features, providing a provable, geometric safeguard during training. Through extensive analysis, we demonstrated that this method sharply blocks the replication of copyrighted details while rigorously retaining semantic fidelity. The gradient projection technique acts fundamentally as a concept acquisition blocker, designed to complement established dememorization safeguards like regularization or data filtering. By integrating our selective learning framework with conventional defenses, we achieved significant reductions in IP leakage risk beyond what those safeguards provide alone. This establishes our approach as a versatile, layer-agnostic addition compatible with any existing training pipeline. The consistent results across different scales validate the scalability and robustness of constraining the gradient update direction. Additionally, the method can be integrated with other post-training techniques for content protection, watermarking [zhao2023recipe, zhang2019robust, xiong2023flexible], fingerprinting [teng2025fingerprinting], and unlearning [wu2025unlearning].
Our method currently achieves robust concept blocking, but the path forward involves maximizing its efficiency and theoretical purity. Our gradient projection approach currently incurs a significant computational cost, nearly doubling training time compared to standard diffusion models due to the need for separate forward and backward passes for copyright-sensitive features. This is a prime area for innovation. Future work can target optimization using low-rank or sparse gradient estimation techniques to dramatically reduce this computational cost. We observed a trade-off with the projection scaling hyperparameter $\lambda$: while the theoretical value of $\lambda=1$ guarantees perfect gradient excising, setting $\lambda=0.1$ works better in practice to ensure gradient stability and optimize convergence speed. This highlights a tension between theoretical completeness and empirical stability. Another direction for future research is to focus on hyperparameter and regularization optimization to maintain training stability while achieving the theoretical guarantee of $\lambda=1$. Further work is needed on refining our approach to successfully disentangle complex copyrighted attributes that remain implicitly entangled with desirable text-conditioned aspects of the image. Additionally, integrating gradient projection into video diffusion pipelines can ensure spatiotemporal copyright compliance without bespoke architectural changes, facilitating cross-modal research and encouraging future advancements in industry best practices for IP protection across the foundation model ecosystem.

### v2

@@ 5.  Conclusion, Limitations, and Future Work
Our current implementation incurs a substantial computational cost, nearly doubling training time relative to standard diffusion models due to the additional forward and backward passes required for copyright-sensitive features. Reducing this overhead is a key direction for future work, potentially through low-rank or sparse gradient estimation. Training stability also requires careful tuning of hyperparameters such as the learning rate $\eta$, the number of training steps, and the projection strength $\lambda$ to prevent divergence. Future research should explore strategies for optimizing $\lambda$ to approach the theoretical guarantee of $\lambda=1$ while maintaining stable learning dynamics. Additional work is needed to disentangle complex copyrighted attributes that remain implicitly coupled with desirable text-conditioned features. Finally, extending gradient projection to video diffusion pipelines offers a promising path toward spatiotemporal copyright compliance without bespoke architectural modifications, enabling cross-modal applications and informing best practices for IP protection across the foundation-model ecosystem.